### v1

# Large Language Model Unlearning for Source Code

Abstract
Large Language Model for Software Engineering (LLM4SE) have demonstrated significant success, but LLMs’ potential memorization of sensitive or outdated training data introduces critical risks to legal compliance, software security, and code quality. LLM unlearning techniques, which can eliminate the influence of undesired data from LLMs in a post-training way, present a promising solution to address these concerns. While recent efforts in LLM unlearning show effectiveness in natural language data, their applicability to source code remains underexplored. Our empirical study reveals that existing LLM unlearning approaches, when applied to source code, cause severe model utility degradation, rendering models practically unusable for code generation. In this paper, we propose PROD, a novel unlearning approach that enables LLMs to forget undesired code content while effectively preserving their code generation capabilities. PROD suppresses the probability of forget data in LLMs’ output distribution while promoting candidate distributional components, enabling the model to jointly learn to forget specific content and retain its general capabilities. To facilitate this study, we establish a benchmark for code unlearning evaluation, which includes three critical downstream tasks: copyrighted code unlearning, insecure code unlearning, and deprecated API unlearning. Our comprehensive evaluation demonstrates that PROD achieves superior balance between forget quality and model utility compared to existing unlearning approaches across three downstream tasks, while consistently exhibiting improvements when applied to LLMs of varying series. PROD also exhibits superior robustness against adversarial attacks without generating or exposing the data to be forgotten. Perceptual quality evaluation further validates our approach’s advantages in both forgetting performance and user experience. These results underscore that our approach not only successfully extends the application boundary of unlearning techniques to source code, but also holds significant implications for advancing reliable code generation.
Index Terms:
Large Language Models, Code Generation, Machine Unlearning
I Introduction
Large language models (LLMs) have revolutionized software engineering, demonstrating remarkable capabilities in various code-related tasks by leveraging training data from massive corpora of open-source code [1, 2, 3]. However, their training data, drawn from broad sources and compiled at massive scale, inevitably contain problematic code, such as copyrighted code files, vulnerable code segments, deprecated APIs, or other undesirable content. With their exceptional memorization capabilities [4], LLMs may retain and reproduce problematic code when assisting users in programming, potentially compromising legal compliance, software security, and code quality. Research [5] has shown that up to 40% of LLM-suggested code may contain security flaws that could lead to serious breaches in production environments. Furthermore, the widespread adoption of LLMs as programming assistants in real-world software development magnifies these concerns, as the issues could lead to significant consequences. Recent high-profile lawsuits against AI companies have drawn considerable attention, with developers challenging AI models for allegedly reproducing proprietary code without permission [6, 7]. This raises a critical question: how can we eliminate the influence of problematic code from LLMs?
Figure 1: Overview of LLM unlearning with the application of removing copyrighted content.
A thorough and direct approach to eliminating the influence of problematic data from LLMs is to retrain LLMs from scratch with clean data. However, this approach incurs prohibitively high computational costs and is impractical for repeated implementation. Recently, LLM unlearning techniques [8, 9, 10, 11, 12] have been proposed to remove the influence of specific training samples while preserving the model’s overall knowledge and functionality, as shown in Figure 1, which holds promise as a potential solution to address the risks posed by problematic data. LLM unlearning techniques leverage post-training strategies to achieve efficient and scalable unlearning without complete model retraining. In natural language processing (NLP) tasks, unlearning techniques have made great progress, such as removing harmful responses, erasing copyrighted content, protecting user privacy, and enforcing policy compliance [9]. However, their application in the domain of code generation remains largely unexplored.
We conduct an empirical study of representative unlearning approaches and their effectiveness in enabling LLMs to forget code. The results show that while these approaches achieve successful forgetting, they often result in significant degradation in model utility. Our analysis reveals that unlearned models often exhibit degraded output, e.g., refusing to generate code, collapsing to high-frequency tokens, producing syntactically incorrect code. We argue that this deterioration stems from the inherent characteristics of code compared to natural language: code exhibits strict syntax, precise dependency modeling, and accurate execution semantics, which are particularly susceptible to disruption. Existing unlearning approaches inadvertently compromise the model’s foundational capabilities in program modeling during the unlearning process. Therefore, It is necessary to develop an unlearning approach applicable to source code to address the problem.
In this paper, we introduce a novel unlearning approach PROD  [^1] PROD stands for Probabilistic Redistribution for Output Distribution of LLMs., which can remove the influence of problematic code from LLMs while effectively preserving the model’s code generation capabilities. PROD optimizes models to fit a target distribution derived from its original output distribution by eliminating the probability of forget data while preserving the distributional structure of the remaining, thereby addressing the model utility degradation of existing unlearning approaches. PROD offers the advantages of lightweight training, efficient optimization, and high usability. We also identify and propose several pressing downstream tasks of LLM unlearning in the domain of source code, including: ❶ copyrighted code unlearning, which enables developers to remove their code contributions from LLMs; ❷ insecure code unlearning, which tackles the issue of LLMs propagating security vulnerabilities; ❸ deprecated API unlearning, which focuses on preventing LLMs from recommending outdated libraries and programming patterns.
Extensive experimental results demonstrate that PROD achieves the best balance between forget quality and model utility among existing LLM unlearning approaches across three downstream tasks. Across four different series of LLMs, our approach consistently demonstrates applicability and effectiveness. To further validate the robustness of PROD, we conduct adversarial attack experiments, and finding that PROD is the only one capable of effectively resisting attacks compared to existing LLM unlearning approaches. Moreover, we conduct perceptual quality evaluation by combining human evaluation and GPT-4 evaluation, and the results shows that PROD has superior forgetting performance coupled with enhanced user experience metrics.
Our main contributions are highlighted as follows:
  • We conduct an investigation of existing LLM unlearning approaches on code-related tasks, revealing that they lead to severe model utility degradation, making models practically unusable for code generation.
  • We propose a novel LLM unlearning approach PROD, that suppresses the probability of forget data in LLMs’ output distribution while promoting candidate distributional components, effectively addressing the model utility degradation of existing unlearning approaches.
  • We identify three critical unlearning applications in the code generation domain (including copyrighted code unlearning, insecure code unlearning, and deprecated API unlearning) and establish a benchmark for evaluating LLM unlearning approaches for code.
  • Extensive experimental results demonstrate that PROD significantly outperforms existing LLM unlearning approaches, achieving successful forgetting of specific code while effectively preserving code generation capabilities.
Our source code and datasets will be released at https://github.com/jiangxxxue/PROD.
II Related Work
II-A Code Generation with LLMs
Code generation is the computational process of automatically producing executable source code from specifications or natural language inputs. Generative models such as ChatGPT [13] and Claude [14] have demonstrated strong capabilities in code generation, benefiting from pretraining on large-scale data and code repositories. Given the immense potential of LLMs, researchers have developed code-specialized LLMs, such as CodeX [15], CodeLlama [16], DeepSeek-Coder [17], Qwen2.5 Coder [18], and CodeGemma [19], which further enhanced LLMs’ capabilities across various software development tasks. However, several critical challenges remain unresolved in leveraging LLMs for code generation, including concerns about safety, legality, and reliability.
Recently, an emerging concern is that LLM-generated code may inadvertently violate security best practices. Because LLMs learn from massive public code repositories, many of which include buggy or outdated patterns, they can propagate known vulnerabilities or poor coding habits. Researchers have begun exploring security-focused code generation techniques [5, 20, 21, 22, 23, 24]. Representative approaches include SaferCode [5], which augments LLM training with carefully curated secure code datasets, and work [20], which proposes incorporating security examples directly into prompts to avoid unsafe coding patterns. Beyond security issues, VersiCode [21] introduces a version-specific benchmark to evaluate how models handle API changes across different releases, revealing that version-controllable code generation poses significant challenges. Failing to adapt can lead to functional errors and long-term maintenance issues. Similarly, SecuCoGen [22] systematically analyzes deprecated API usage in LLM-based code completion. Testing seven advanced models on over 28k prompts and 145 API changes, they observe a substantial rate of outdated suggestions, indicating that many LLMs fail to migrate to newer APIs. Additionally, copyright concerns have also been raised regarding LLM-generated code. LiCoEval [23] systematically evaluates LLMs’ license compliance and finds that models sometimes produce code fragments that closely resemble copyrighted implementations, often without proper attribution.
These limitations highlight that despite remarkable progress, LLM-based code generation still falls short of developers’ expectations for producing secure, legally compliant, and up-to-date code. Our work aims to address these gaps by proposing an effective unlearning approach for suppressing undesired code output from LLMs.
II-B LLM Unlearning
Machine unlearning, the process of removing specific data and its influence from a trained machine learning model, has garnered significant attention in recent years. Since [25] first introduced the concept of machine unlearning, laying the groundwork for subsequent research. Due to the widespread application and the relative simplicity of classification models, they become the focal point for most of the machine unlearning works [26, 27, 28, 29, 30, 31]. Machine unlearning approaches can be broadly categorized into three types: data-reversed training [32, 33, 34], influence function based approaches [28], and optimization based unlearning [35, 36]. Data-reversed training techniques employ noise injection to reverse or weaken the model’s memory and dependence on specific data. Influence function-based approaches evaluate the impact of training samples on the model by perturbing model parameters, which are less common in LLM unlearning due to the computational complexity of Hessian matrix inversion. Optimization-based unlearning, which is the focus of our work, achieves forgetting through direct model optimization. The rapid advancement of LLMs presents more significant challenges for machine unlearning. Their massive parameter scales and generative nature make traditional unlearning approaches unsuitable [8]. Recent research efforts have been dedicated to the domain of LLM unlearning within the field of NLP, proposing methods such as GA [9], DPO [11], NPO [10], and FLAT [12], which aim to address forgetting tasks in the field of NLP, such as user privacy unlearning, copyrighted books unlearning, and harmful content unlearning [9]. To facilitate the evaluation of LLM unlearning, several benchmarks have been proposed. These include synthetic profiles and question-answer pairs about fictitious authors [37], as well as the forgetting of specific content like Harry Potter books [38]. Moreover, to assess the vulnerabilities of unlearned models, various attack approaches have been employed, such as prefix injections [39, 40, 41]. These attacks help researchers understand the robustness of unlearning techniques and identify areas for improvement.
Given that source code possesses unique characteristics compared to natural language, although the exploration of LLM unlearning in NLP has proven its value, its application in code generation remains largely unexplored. This necessitates the development of LLM unlearning approaches applicable to code to address the challenges.
III Preliminary Investigation on LLM Unlearning
We first give the definition of LLM unlearning problem. Then we introduce four representative LLM unlearning approaches. Finally, we investigate the performance of existing LLM unlearning approaches on code-related tasks and elaborate on the severe degradation issues in model utility caused by existing unlearning approaches when applied to source code.
III-A Problem Definition
For the unlearning problem, we begin with forget data $D_{f}$ and a LLM $\pi_{\theta}$ that needs to undergo unlearning. $D_{f}$ represents the content we wish the model to forget. In practical scenarios, $D_{f}$ might be a collection of model-generated code that has been reported by users for copyright infringement. $D_{f}$ consists of inputs $x_{f}$ and outputs $y_{f}$, where the inputs represent prompts that might trigger undesired outputs $y_{f}$ [^2] Inputs may be omitted in some scenarios, such as when the goal is to forget entire code content, the prompt can be empty.. $D_{f}$ does not need to be an exact subset of $\pi_{\theta}$’s training data. The objective of LLM unlearning is to efficiently remove the impact of $D_{f}$ on $\pi_{\theta}$, meaning that the unlearned model $\pi_{\theta}^{*}$’s outputs should be as distant as possible from $y_{f}$.
We evaluate the effectiveness of an LLM unlearning approach from two dimensions: forget quality and model utility. Forget Quality refers to the degree to which the unlearned model eliminates the influence of target forget data, while Model Utility refers to the general performance of the model on other tasks. An ideal unlearning approach aims to increase forget quality while minimizing the reduction in model utility.
III-B LLM Unlearning Approaches
Gradient Ascent (GA) [9]. As a simple and widely used unlearning approach, GA [9] performs gradient ascent on next-token prediction loss over the target forget data to approximately ”reverse” the effects of gradient descent optimization that occurred during training on the forget data, formally denoted as $L_{GA}$.
$\mathcal{L}_{\text{GA}}(\theta)=-\underbrace{\mathbb{E}_{\mathcal{D}_{f}}[-\log(\pi_{\theta}(y_{f}|x_{f}))]}_{\text{prediction loss}}=\mathbb{E}_{\mathcal{D}_{f}}[\log(\pi_{\theta}(y_{f}|x_{f}))].$ (1)
However, GA suffers from divergent properties and excessively rapid linear convergence rates, resulting in unstable training dynamics and potential model collapse. NPO is developed to address the limitation. NPO is inspired by DPO, which is a well-established approach in the field of LLM preference optimization.
Direct Preference Optimization (DPO) [11]. DPO enables the model to learn to increase the probability of preferred responses $y_{w}$ while decreasing the probability of non-preferred responses $y_{l}$, all while optimizing under the constraint of staying close to a reference model. The loss function $\mathcal{L}_{\text{DPO}}(\theta)$ is defined as:
$\footnotesize{\mathcal{L}_{\text{DPO}}(\theta)=-\mathbb{E}_{(x,y_{w},y_{l})}\left[\log\sigma\left(\beta\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right)\right],}$ (2)
where $(x,y_{w},y_{l})$ represents the training data triplets: prompt, preferred response, and non-preferred response, $\pi_{\theta}$ is the model being trained, $\pi_{\text{ref}}$ is the reference model (typically the initial LLMs for training), $\beta$ is a hyperparameter that controls the degree of deviation from the reference model and $\sigma$ is the sigmoid function.
Although DPO is originally designed for LLM preference optimization, researchers have discovered that it can be applied to unlearning tasks by providing template responses $y_{e}$ (such as ”I don’t know”) as the preferred response $y_{w}$, and $y_{f}$ as non-preferred response $y_{l}$.
Negative Preference Optimization (NPO) [10]. Since unlearning tasks typically only provide negative samples (i.e., forget data), NPO retains only the log-probability ratio for non-preferred responses from the DPO formulation to steer the model away from generating undesired content. The NPO loss function $\mathcal{L}_{\text{NPO}}(\theta)$ is defined as:
$\mathcal{L}_{\text{NPO}}(\theta)=-\mathbb{E}_{(x_{f},y_{f})}\left[\log\sigma\left(\beta\left(-\log\frac{\pi_{\theta}(y_{f}|x_{f})}{\pi_{\text{ref}}(y_{f}|x_{f})}\right)\right)\right],$ (3)
where NPO leverages the structural advantages of the DPO loss formulation (specifically the $\log\sigma$ transformation and the regularization provided by $\pi_{\text{ref}}$) to achieve logarithmic convergence rates, which typically yields more stable training dynamics compared to GA approach.
Forget Data Only Loss Adjustment (FLAT) [12]. FLAT proposes to use a loss function that balances two key objectives: discouraging the generation of undesirable responses to forget data $y_{f}$ while simultaneously promoting the production of appropriate template responses $y_{e}$ when encountering relevant forget prompts,
$\mathcal{L}_{\text{FLAT}}(\theta)=-\lambda_{f}\cdot L_{f}(x_{f},y_{f};\theta)+\lambda_{e}\cdot L_{e}(x_{f},y_{e};\theta),$ (4)
where $\lambda_{f}$ and $\lambda_{e}$ are wight to balance between $L_{e}(x_{f},y_{e};\theta)$ and $L_{f}(x_{f},y_{f};\theta)$. FLAT introduces f-divergence to establish an appropriate equilibrium between the two objectives. The loss function is subsequently reformulated into a variational form of f-divergence:
$\mathcal{L}^{\text{FD}}_{\text{FLAT}}(\theta)$ $=-\mathbb{E}_{(x_{f},y_{e},y_{f})}\left[g^{*}(\mathbb{P}(x_{f},y_{e};\theta))+f^{*}(g^{*}(\mathbb{P}(x_{f},y_{f};\theta)))\right],$ (5)
$\mathbb{P}(x_{f},y;\theta)=\frac{1}{|y|}\sum_{t=1}^{|y|}\frac{\pi_{\theta}(y^{1:t}\mid x_{f})}{\pi_{\theta}(y^{1:t-1}\mid x_{f})},$
where $y$ represents the template answer $y_{e}$ or $y_{f}$, and $y^{1:t}$ denotes the first $t$ tokens of the sequence $y$. FLAT automatically assigns appropriate weights $g^{*}$ and $f^{*}(g^{*}())$ to the loss components. f-divergence is a divergence framework that includes various divergence measures, such as the Kullback-Leibler (KL) divergence, Jensen-Shannon (JS) divergence, and Pearson divergence. $g^{*}$ and $f^{*}(g^{*}())$ can be configured according to the specific divergence measure chosen.
Figure 2: Performance of LLM unlearning approaches on source code. The upper and lower subplots respectively show Forget Quality and Model Utility across training epochs. Forget quality reflects how well the model forgets target data, while model utility indicates the retention of general capabilities. For both metrics, higher values indicate better performance.
Figure 3: Cases of different LLM unlearning approaches on forget quality and model utility.
III-C Empirical Study of Code Unlearning
We investigate the effectiveness of existing LLM unlearning approaches for source code. Taking the copyrighted code unlearning task as an example, we evaluate both the forget quality on copyrighted code and the model utility on code generation. In copyrighted code unlearning task, it requires that after applying unlearning approaches, the LLM previously exposed to copyrighted code no longer reproduces such content, even when prompted with inputs resembling copyrighted contexts. [^3] Detailed benchmark descriptions and experimental settings are provided in Section V-A and Section V-B.
As shown in Figure 2, our experimental results reveal that although these approaches demonstrate satisfactory forget quality (approximately 90%), they perform poorly in maintaining model utility (reduction exceeding 60%). As training progresses, model utility decreases dramatically, eventually reaching an unacceptable level despite achieving the desired forget quality. This indicates that although the LLM successfully forgets the target code, it concurrently loses essential coding capability, rendering it practically unusable, which is undesirable. Our goal is to avoid generating target code while retaining a usable LLM.
[…]
Our preliminary findings reveal that source code possesses unique characteristics and challenges compared to natural language data, which existing LLM unlearning approaches struggle to handle these differences. This necessitates the development of advanced unlearning approaches to address the challenge.
IV Methodology
Figure 4: The PROD unlearning pipeline. Capital letters represent the corresponding section numbers in methodology.
We introduce PROD as a novel LLM unlearning approach, which can remove the influence of unwanted code data from LLMs while significantly mitigating the model utility degradation problem of previous LLM unlearning approaches on code-related tasks. PROD is not only robust but also training-efficient, positioning it as a feasible and practical option for applications in code LLMs.
Motivation. Existing LLM unlearning approaches typically involve designing specialized loss functions to optimize the model. However, their failure can be mainly attributed to their lack of mechanisms to preserve the model’s original capabilities, leading to the catastrophic forgetting of all relevant knowledge. For tasks such as code generation, certain foundational knowledge (e.g., basic syntax, core logic) is indispensable, and its unlearning would lead to a complete breakdown of model’s functionality. Therefore, establishing a mechanism to retain core competencies is imperative and serves as the driving force behind our work.
Overview. PROD achieves targeted unlearning by directly manipulating the LLM’s output probability distribution. The core principle is to suppress the generation probability of unwanted tokens while maintaining the distribution structure over non-targeted tokens, thereby preserving general model capabilities. The methodology involves three key stages. First, for data instances designated for unlearning, the model’s original output probability distributions are collected (Section IV-A). Second, an “ideal” target distribution is constructed. This involves identifying probability components in the original distribution that correspond to the content to be unlearned. The probabilities of these target tokens are then nullified, and their mass is redistributed among the remaining tokens in a manner that preserves their relative likelihoods (Section IV-B-IV-D). Third, the LLM is fine-tuned using a specialized loss function. This optimization process guides the model’s output distribution to converge towards the constructed ideal target distribution, thereby effecting the unlearning of the specific content (Section IV-E).
In the following sections, we will present a detailed description of PROD, with its illustration shown in Figure 4.
IV-A Output Distribution Collection.
The initial phase of PROD involves acquiring the original output probability distributions from LLM for data instances designated for unlearning. For the forget data $D_{f}=\{(x_{f},y_{f})\}$, each $x_{f}$ is an input prompt, and $y_{f}=(y_{f,1},y_{f,2},\dots,y_{f,L})$ is the corresponding output sequence targeted for removal. For each instance $(x_{f},y_{f})\in D_{f}$, and for each token $y_{f,t}$ in the sequence $y_{f}$ (where $t$ ranges from $1$ to $L$), we obtain the LLM’s prediction. Specifically, given the input prompt $x_{f}$ and the preceding ground-truth tokens $y_{f,<t}=(y_{f,1},\dots,y_{f,t-1})$, the model produces a logit vector $h_{t}\in\mathbb{R}^{|V|}$, where $V$ denotes the model’s vocabulary, and $|V|$ is its size. These logits $h_{t}$ are subsequently transformed into an original probability distribution $p(\cdot|x_{f},y_{f,<t})$ using the softmax function. This distribution $p$ quantifies the model’s initial likelihood for each token $w^{i}\in V$ to be the next token, conditioned on the prompt $x_{f}$ and the prefix $y_{f,<t}$:
$p(w^{i}|x_{f},y_{f,<t})=\text{Softmax}(h_{t})^{i}=\frac{\exp(h_{t}^{i})}{\sum_{k=1}^{|V|}\exp(h_{t}^{k})},$ (6)
This collection process yields the raw distributions that will be modified in subsequent steps of PROD.
IV-B Forget Data Elimination.
To unlearn a target sequence $y_{f}$ for an input $x_{f}$, we iteratively modify the output probability distribution at each token generation step $t$. The principle is to nullify the probability of the target token $y_{f,t}$ (given $x_{f},y_{f,<t}$) and reallocate its probability mass. For each target token $y_{f,t}$, we operate on its corresponding logit vector $h_{t}$ (from Eq. 6). A binary mask $m_{t}\in\{0,1\}^{|V|}$ is defined where $m_{t}^{i}=1$ if vocabulary token $w^{i}=y_{f,t}$, and 00 otherwise. This mask transforms $h_{t}$ into $\hat{h}_{t}$ by setting the logit for $y_{f,t}$ to $\infty$- ∞:
$\hat{h}_{t}^{i}=\begin{cases}-\infty,&\text{if }m_{t}^{i}=1,\\ h_{t}^{i},&\text{otherwise},\end{cases}$ (7)
where $j\in\{1,\dots,|V|\}$. The modified logits $\hat{h}_{t}$ are then normalized using softmax to yield a new distribution $p_{n}(\cdot|x_{f},y_{f,<t})$:
$p_{n}(w^{i}|x_{f},y_{f,<t})=\text{Softmax}(\hat{h}_{t})^{i}=\frac{\exp(\hat{h}_{t}^{i})}{\sum_{k=1}^{|V|}\exp(\hat{h}_{t}^{k})}.$ (8)
This ensures $p_{n}(y_{f,t}|x_{f},y_{f,<t})$ is effectively zero. The probability mass of $y_{f,t}$ is reallocated proportionally among other tokens in $V\setminus\{y_{f,t}\}$, preserving their relative probabilities.
IV-C Noise Elimination.
Suppressing target token probabilities (Forget Data Elimination) can inflate probabilities of non-target, low-probability tokens, introducing noise into $p_{n}(\cdot|x_{f},y_{f,<t})$. To mitigate this, Nucleus Sampling [^4] Nucleus Sampling involves choosing from a smaller set of plausible candidates by dynamically selecting a variable-sized subset of tokens that cumulatively make up a certain probability mass (e.g., top 80%). truncates the distribution’s tail. The logits $\hat{h}_{t}$ are transformed into $\widetilde{h}_{t}$:
$\widetilde{h}^{i}_{t}=\begin{cases}-\infty,&\text{if }w^{i}\notin S_{p},\\ \hat{h}^{i}_{t},&\text{otherwise,}\end{cases}$ (9)
where $S_{p}$ is the smallest set of tokens $w^{i}$, sorted by descending $p_{n}(w^{i}|x_{f},y_{f,<t})$, whose cumulative probability is at least $p$. We empirically set the nucleus threshold $top\_p=0.8$. Applying Forget Data Elimination before Noise Elimination is critical. Reversing this order could retain target tokens or result in an overly sparse distribution. This sequence ensures effective unlearning while maintaining a robust distribution over remaining tokens.
IV-D Probabilistic Redistribution.
Finally, the target probability distribution $p_{\text{T}}$ (supervisory signal) is defined. It adjusts probabilities from noise-eliminated logits $\widetilde{h}_{t}$, penalizing the original probability of the token to be unlearned. For each vocabulary token $w^{i}$, given $x_{f}$ and prefix $y_{f,<t}$, $p_{\text{T}}(w^{i}|x_{f},y_{f,<t})$ is:
$p_{\text{T}}(w^{i}|x_{f},y_{f,<t})$ $=\text{Softmax}(\widetilde{h}_{t})^{i}-\alpha\cdot\text{Softmax}(h_{t})^{i}\cdot m_{t}^{i}$ (10)
=$=\begin{cases}\text{Softmax}(\widetilde{h}_{t})^{i},&\text{if }m_{t}^{i}=0,\\ -\alpha\cdot p(w^{i}|x_{f},y_{f,<t}),&\text{otherwise,}\end{cases}$
where $\alpha$ is another hyperparameter controlling the enhanced suppression degree of forget data in each step, which is set to 0 by default. In practice, we find that as gradients update during training, the model’s output distribution changes. If the target distribution built on the distribution also constantly varies, the trained model’s output tends to collapse to the most frequently occurring token. Therefore, we consistently use the target distribution constructed from the initial model’s distribution throughout training to ensure stability.
IV-E Loss Function.
Through these operations, we construct the final target distribution $p_{\text{T}}$ as the supervisory signal for training. We then select appropriate loss functions to optimize the model’s output to fit the target distribution, i.e., the cross-entropy loss:
[…]
where $T$ is the sequence length and $p(y_{t}^{i}|x_{f},y_{<t})=\text{Softmax}(h_{t})$. The cross-entropy loss function offers several advantages in our context, which can naturally accommodate target distributions containing negative probability values that arise from our distribution transformation procedure. This property is essential for effectively implementing the suppression of undesired responses while promoting alternative, acceptable outputs. While cross-entropy loss serves as our primary loss function, it is worth noting that alternative distribution distance measures could also be employed within this framework, e.g., Kullback-Leibler (KL) divergence or Jensen-Shannon (JS) divergence could be utilized to quantify the discrepancy between the model’s output distribution and our target distribution.
During the optimization process, we update the model parameters using gradient descent according to the loss:
[…]
Algorithm 1 PROD Unlearning Procedure
0:  LLM $\pi_{\theta}$; Forget set $\mathcal{D}_{f}=\{(x_{f},y_{f})\}$; Learning rate $\eta$; Training Epochs $N$; Hyperparameters $top\_p,\alpha$.
0:  Unlearned LLM $\pi_{\theta}^{*}$.
1:  for k $=0$ to $N-1$ do
2:     for each $(x_{f},y_{f})\in\mathcal{D}_{f}$ do
3:        Collect output distribution $p(w|x_{f},y_{f,<t})$ via Eq. (6).
4:        Compute $\hat{h}_{t}$ and $p_{n}(w|x_{f},y_{f,<t})$ to apply forget data elimination via Eq. (7) and Eq. (8), respectively.
5:        Compute $\widetilde{h}_{t}$ to perform noise elimination via Eq. (9).
6:        Compute probabilistic redistribution $p_{\text{T}}(w|x_{f},y_{f,<t})$ based on $\widetilde{h}_{t}$, $\alpha$, and $p(w|x_{f},y_{f,<t})$ via Eq. (10).
[…]
9:        $k\leftarrow k+1$ and $\theta^{*}\leftarrow\theta_{k}$
10:     end for
11:  end for
12:  return  $\pi_{\theta}^{*}$
In summary, the pipeline of PROD is outlined in Algorithm 1 and PROD presents two advantages: 1) Model Utility Preservation. By constructing the target distribution through controlled modification of the original distribution, we maintain the model’s fundamental capabilities while selectively removing specific content; 2) Optimization Efficiency. The proximity between our constructed target distribution and the model’s original output distribution facilitates more efficient convergence during the optimization process.
V Evaluation
PROD aims to remove the impact of problematic code data (e.g., copyrighted code files, vulnerable code segments, deprecated APIs) from LLM, addressing the problem of severe model utility degradation observed in existing LLM unlearning approaches. In this section, we evaluate the effectiveness of our approach PROD through five research questions:
  • RQ1: How does PROD perform on forget quality and model utility across different downstream tasks?
  • RQ2: How well does PROD apply to LLMs across diverse series?
  • RQ3: How robust is PROD against adversarial attacks?
  • RQ4: How does each component of PROD contribute to its overall effectiveness?
  • RQ5: How does PROD perform in forgetting effectiveness from a perceptual quality perspective?
Through these research questions, we provide a comprehensive evaluation of PROD to examine performance, applicability, unlearning robustness, component contributions, and user experience in the context of unlearning for source code.
V-A Code Unlearning Benchmark
Since current benchmarks for evaluating LLM unlearning approaches primarily focus on NLP tasks, we need to develop an evaluation benchmark specifically for code generation. By constructing data, or reusing existing datasets and designing evaluation metrics, we establish a code unlearning benchmark to evaluate the performance of unlearning approaches for source code. Our benchmark focuses on three key downstream tasks: copyrighted code unlearning, insecure code unlearning, and deprecated API unlearning. These tasks are designed to evaluate forget quality. Additionally, we incorporate a general code generation dataset to evaluate model utility.
Copyrighted Code Unlearning. We randomly sample and select 100 code files from the pretraining dataset The Stack [42] as the forget set $D_{FG}$ to simulate scenarios where users want LLMs to forget private code assets. The Stack is a high-quality code corpus that has undergone rigorous filtering and deduplication, widely used for model training. We manually review the sampled code files, eliminating test files, simple examples, and framework templates with low information content. Our selection criteria require each code file to contain meaningful functionality implementations, business logic, or algorithm designs, thereby reflecting intellectual property code that needs protection in real-world scenarios. The objective of this task is to ensure that the unlearned model does not generate code with high similarity to the copyrighted code files. During unlearning, we forget the entire code file. Therefore, the Copyrighted Code Unlearning task has the longest code length to be forgotten, making it the most difficult unlearning task among the three downstream tasks. When evaluating the forget quality metric, we provide the first half of a code file (determined by splitting the file using tokenizer) as a prompt to the model. Given that copyright law protects expression rather than ideas [43, 44], we employ textual similarity metrics to assess potential copyright infringement in model outputs. Therefore, the forget quality is measured by the complement of the BLEU score [45] between the model’s output and the actual latter half of the code file, where a higher score indicates better forgetting of the copyrighted code.
Insecure Code Unlearning. We establish the insecure code unlearning task using the CyberSecEval dataset from Purple Llama Project, which is designed to assess the cyber security vulnerabilities of LLMs. CyberSecEval has collected instances of insecure coding practices from open-source code, which we utilize as forget data. The dataset contains 1,916 samples covering 8 programming languages and 50 Common Weakness Enumeration (CWE) defined insecure coding practices. Forget quality is evaluated in autocomplete contexts, that is given the code before the vulnerable code segment (where vulnerability locations have been identified), we prompt the model to generate subsequent code and observe whether the generated code contains vulnerabilities. The forget quality metric is computed as the average of two distinct measures. The first measure is the complement of the BLEU score calculated between the model’s output and the reference code. The second measure is the pass rate obtained through CyberSecEval’s accompanying static analysis tool, which detects insecure code lines.
(a) Copyrighted Code Unlearning
(b) Insecure Code Unlearning
(c) Deprecated API Unlearning
Figure 5: Forget quality versus model utility across different unlearning tasks, where the length of forget code snippets gradually decreases from (a) to (c). Performance curves positioned closer to the upper-right corner indicate superior approach effectiveness.
Deprecated API Unlearning. We construct the deprecated API unlearning task based on a version-controllable code generation dataset (Versicode) [21]. To adapt the original Versicode dataset for our deprecated API unlearning purpose, we perform several processing steps. We extract the meta data (i.e., package names and version numbers) from Versicode, retaining only packages with explicit version declarations (e.g., ==2.1.0) while filtering out those with ambiguous version constraints (e.g., >=) or without version annotations. Additionally, we only keep packages with at least 3 data samples. After data cleaning, we divide the APIs into deprecated and still-valid categories by selecting a middle version number as the split point for each package, consequently dividing all data samples into two parts. The final processed dataset contains 252 packages, with an average of 14 samples and 10 different APIs per package, totaling 3449 data samples. For this task, we leverage all deprecated API data to perform unlearning while checking whether the unlearned model can generate valid APIs on the remaining samples. In the prompt, we specify a valid API version number to guide the model’s output. The task is formulated as code completion, where given the context before an API, the model predicts the API, and we use exactly match metric as forget quality.
To evaluate model utility, following previous work [5], we employ HumanEval [15], a widely adopted benchmark for general code generation. For each programming problem in the benchmark, we measure the PassRate of the generated code by executing test cases. For each task, $n\geq 1$ samples of code are generated, and the number of samples, $c\leq n$, that pass the test cases are counted. The model utility is computed as the average PassRate on these code generation datasets, where PassRate is defined as: $\operatorname{PassRate}=\mathop{\mathbb{E}}\limits_{\operatorname{Problems}}\begin{bmatrix}1-\tiny{\frac{\begin{pmatrix}n-c\\ 1\end{pmatrix}}{\begin{pmatrix}n\\ 1\end{pmatrix}}}\end{bmatrix}.$
Due to the inherent trade-off between forget quality and model utility, evaluating methods via simple averaging is ineffective. To address this, we introduce the Pareto Dominance Ratio (PDR), measuring a method’s relative superiority in multi-objective scenarios. Given approach $\mathcal{M}={m_{1},\dots,m_{n}}$ and objectives $\mathcal{O}$, PDR is defined as:
$\text{PDR}(m_{i})=\frac{|\{m_{j}\mid m_{i}\succ m_{j},m_{j}\in\mathcal{M},m_{j}\neq m_{i}\}|}{|\mathcal{M}|-1},$
where $m_{i}\succ m_{j}$ indicates $m_{i}$ is no worse than $m_{j}$ across all objectives and strictly better in at least one objective.
V-B Implementation Details
We detail the specific implementation details of our evaluation from three aspects. Specifically,
Configurations for Downstream Tasks. For copyrighted code unlearning and insecure code unlearning tasks, following previous unlearning work [9, 10, 12], we first train the model on forget data $D_{FG}$ to ensure the model has memorized the content before applying unlearning approaches. We use a learning rate of 2e-5, batch size of 32, and train for 10 epochs. To prevent overfitting on $D_{FG}$ during training, we employ continued pretraining to train the model, where in each epoch, the forget data is mixed with different pretraining data in a ratio of approximately 1:9, which is a subset of The Stack corpus. For deprecated API unlearning task, we directly apply LLM unlearning approaches on the LLMs without training on forget data. This is because the APIs we target in this task are common functions from widely used libraries that LLMs have basically encountered during pretraining.
Training Setup for Unlearning. We use CodeLlama-7B [16] as our base model by default for our evaluation. All experiments are conducted on 4 NVIDIA A100 GPUs. Following previous work [10], we set the batch size to 32 and train for 10 epochs. For all unlearning approaches, we select the learning rate via grid search from the set {1e-4, 5e-5, 1e-5, 5e-6, 1e-6}, optimizing for the forget quality metric on forget data. We use AdamW optimizer [46] with a weight decay of 0.01 and set the maximum sequence length to 1024 tokens during training. During training, we only optimize the loss over the forgetting part of each example. The unlearning training is conducted five times using different random seeds, and the final evaluation results are averaged across the five runs.
Configurations for Baselines and PROD. NPO and DPO use the hyperparameter $\beta$ of 0.1. FLAT employs KL divergence in its loss, which demonstrated superior performance in the original study, where $g^{*}(v)=v$ and $f^{*}(u)=e^{u-1}$. For our approach PROD, we set the default values of $top\_p$ and $\alpha$ to 0.8 and 0, respectively.
V-C Evaluation Results
In this section, we present the experimental results addressing our research questions and highlight key findings that emerged from our comprehensive evaluation.
V-C1 RQ1: Performance Comparison
To evaluate the effectiveness of PROD on code unlearning, we compare PROD with existing representative unlearning approaches by assessing both forget quality and model utility across three downstream tasks: copyrighted code unlearning, insecure code unlearning, and deprecated API unlearning.
Setup. In this evaluation, we assess forget quality and model utility across various training epochs for different machine unlearning approaches. During inference, we employ greedy decoding (temperature = 0) for generation. We define the contaminated model as the original model, and the model after applying unlearning approaches as the unlearned model. The performance of original model serves as the baseline.
TABLE I: PDR comparison between different unlearning approaches, where higher PDR indicates greater effectiveness, demonstrating that the evaluated approach consistently outperforms other approaches across multiple objectives.
Task GA DPO NPO FLAT Ours
Copyrighted Code Unlearning 0.153 0.036 0.006 0.000 0.418
Insecure Code Unlearning 0.133 0.000 0.245 0.061 0.704
Deprecated API Unlearning 0.315 0.206 0.516 0.000 0.580
Results. Figure 5 shows the performance of forget quality and model utility for various unlearning approaches in three downstream tasks. From the overall performance comparison, our approach demonstrates a consistent trend towards the upper-right corner compared to other approaches, indicating a superior balance between forget quality and model utility. The PDR metric quantifies this phenomenon. As shown in Table I, our approach achieves the highest PDR across all tasks, demonstrating a significant average relative improvement of 124% compared to the best-performing baseline. During training, our approach exhibits rapid forgetting in the early stages, followed by gradual convergence in both forget quality and model utility. This pattern reflects high forgetting efficiency while effectively maintaining model utility. These advantages stem from our approach’s underlying computational mechanism, which fits a target distribution and employs gradient descent optimization, eventually converging to a bound and effectively preventing model collapse. Moreover, we observe varying degrees of model utility decline across different downstream tasks, resulting in performance curves with different slopes. Steeper slopes indicate more severe utility degradation. Among these, Copyrighted code unlearning task is the most difficult in maintaining utility, which we find positively correlates with the length of content to be forgotten. Notably, PROD demonstrates significant advantages even in this most difficult task.
V-C2 RQ2: Application on Different LLMs
To investigate the broad applicability of our approach, we evaluate its performance across a diverse range of LLMs.
Setup. We implement our unlearning approach on four representative code LLMs, including CodeLlama-7B [16], Qwen2.5-Coder-7B [18], Deepseek-coder-6.7B [17], and Starcoder-7B [47], which differ in their architectural designs and training recipes. To ensure experimental rigor, we maintain consistency across all other settings while observing performance variations across training epochs on the copyrighted code unlearning task.
Figure 6: The performance of PROD on different LLMs, where the dotted line represents the best model utility of all other approaches in achieving comparable (i.e., 90%) forget quality.
Results. The experimental results are illustrated in Figure 6. The results shows that PROD has similar performance trends across all four LLMs: as training progresses, all LLMs achieve near-perfect forget quality while maintaining stability and converging in terms of model utility, which significantly outperforms all baseline approaches in model utility, highlighting the superiority and robustness of PROD. Moreover, CodeLlama-7b and Starcoder-7b demonstrate better model utility preservation compared to other LLMs, suggesting that the effectiveness of unlearning may be influenced by the inherent stability of the underlying LLMs.
V-C3 RQ3: Adversarial Attacks
Recent studies have revealed the vulnerability of LLM unlearning approaches [40, 41], where forgotten content resurfaces under certain adversarial attacks. Therefore, we employ classical prefix injection attacks to evaluate the robustness of PROD against adversarial perturbations, using various prefix lengths.
Setup. We implement prefix injection attacks [40, 41] by providing a prefix of forgotten content and observing whether unlearned models obtained through different LLM unlearning approaches generate the subsequent forgotten content. By calculating the BLEU similarity score between the model’s response and the original copyrighted code, we can quantify the effect of the attack: higher BLEU scores indicate that the model can still regenerate the forgotten content, suggesting insufficient robustness of the unlearning approach. Moreover, to ensure meaningful evaluation, we only test unlearned model checkpoints that retain at least 60% of the original model’s performance on code generation. This criterion helps exclude models that appear resistant to attacks merely due to degraded functionality rather than successful unlearning.
Figure 7: Comparison of adversarial attack results across different LLM unlearning approaches, showing mean attack effects with maximum and minimum ranges.
Results. As shown in Figure 7, the experimental results of adversarial attacks demonstrate that PROD significantly outperforms all other LLM unlearning approaches. Specifically, when subjected to prefix injection attacks, PROD maintains a notably low similarity (approximately 0.05) between generated content and forgotten content, while other LLM unlearning approaches exhibit substantially higher similarity scores (all exceeding 0.3). Furthermore, PROD shows the smallest variance, indicating strong stability in performance. These two aspects provide compelling evidence that compared to existing LLM unlearning approaches, our approach achieves superior robustness in unlearned model.
V-C4 RQ4: Ablation Study
To analyze the effectiveness of our approach, we conduct a comprehensive ablation study. This evaluation includes using alternative loss and adjusting critical hyperparameter configurations.
(a) Loss Function
(b) Hyperparameter $\mathbf{top\_p}$
(c) Hyperparameter $\mathbf{\alpha}$
Figure 8: Ablation results on alternative loss function, and the impact of hyperparameters $top\_p$ and $\alpha$ in PROD.
Setup. We replace the cross-entropy loss with two widely adopted distribution divergence loss, i.e., KL divergence loss and JS divergence loss. We also investigate the impact of hyperparameters $top\_p$ and $\alpha$. Specifically, $top\_p$ controls the degree of noise elimination by filtering out low-probability portions in the output distribution, and $\alpha$ regulates the enhanced suppression degree of forget data in each update step.
Results. The results of ablation study are presented in Figure 8. From Figure 8(a), we observe that cross-entropy loss demonstrates better performance compared to KL divergence and JS divergence. Figure 8(b) demonstrates that noise elimination contributes to improved unlearning performance, where moderate noise removal ($top\_p=0.8$) achieves optimal results and no noise removal ($top\_p=1$) yields the worst performance. Although excessive noise removal ($top\_p=0.2$) shows some improvement, it fails to match the effectiveness of the moderate setting. Figure 8(c) shows that the forgetting regulation parameter $\alpha$ has a large impact on unlearning performance. When $\alpha=0$, which prohibits the generation of forget data by setting its probability to zero in the target distribution, the model achieves optimal performance. Positive $\alpha$ values, although theoretically enhancing forgetting through gradient ascent, practically lead to model collapse, with larger values accelerating this deterioration. Small negative $\alpha$ values, while achieving some degree of forgetting, fail to match the effectiveness of complete prohibition ($\alpha=0$).
V-C5 RQ5: Perceptual Quality Evaluation
Following the work [48, 49], we conduct a perceptual quality evaluation to assess the subjective human experience of our approach.
Setup. We invite three independent volunteer evaluators (each with two years or more of software development experience). We randomly select 20 samples from each of the three downstream tasks, including generation results from different approaches. The evaluation is conducted in a comparative manner, where evaluators are presented with samples from our approach and baseline approaches, and asked to select the better one with a justification. Each human evaluator assesses all samples without knowing which approach produced each sample. We calculate the win rate of our approach against baselines. The win rate is calculated as the total number of times our approach outperformed baselines divided by the total number of comparisons in the evaluation.
Given that human evaluation can only assess a limited portion of samples due to practical constraints, and recent research suggests that LLMs can substitute human evaluation to some extent with high evaluation consistency [50], we introduced GPT-4 to evaluate all samples across all downstream tasks.
TABLE II: Win rate comparison on perceptual evaluation.
Comparison Human Evaluation GPT-4 Evaluation
Ours vs GA 81% 77%
Ours vs DPO 92% 86%
Ours vs NPO 76% 73%
Ours vs FLAT 87% 81%
Results. The results of the perceptual quality evaluation are shown in Figure II. We find that our approach achieves win rates above 70% compared to baselines. According to the reasons provided by human evaluators, the common feedback was that our approach still produced valuable outputs after forgetting specific content, while baselines often produced empty outputs or messy code.
VI Threats to Validity
We have taken all reasonable steps to mitigate potential threats that could hamper the validity of this work. Specifically, we identify and attempt to mitigate two main categories of threats: internal validity threats and external validity threats. Internal threats to validity pertain to factors that may impact the observed causal relationships in our experiments. A potential threat arises from the sensitivity of LLM unlearning approaches to hyperparameter settings, especially learning rates. To mitigate this, we conduct extensive hyperparameter tuning through grid search for learning rates and report by averaging multiple results. We kept other hyperparameters consistent with established practices from prior studies to minimize unintended biases. External threats to validity concern the generalizability of our findings beyond the experimental contexts considered in this study. We address this by evaluating PROD on three representative unlearning tasks, each reflecting distinct practical scenarios encountered in real-world software engineering contexts. Additionally, we validate our approach across multiple LLM architectures, including CodeLlama, Deepseek-Coder, Qwen2.5-Coder, and Starcoder, which vary in their training data, model architectures, and parameter scales. Thus, we believe the concern about the generalizability of our findings is mitigated.
VII Conclusion
In this paper, we have investigated existing LLM unlearning approaches on source code, identifying significant utility degradation. To this end, we propose PROD, a novel unlearning approach that reconstructs the output distribution by suppressing target code and preserving essential code-generation capabilities. We also establish a benchmark covering copyrighted code, insecure code, and deprecated API unlearning tasks. Extensive experiments show PROD significantly outperforms existing methods in both forgetting performance and user experience, while also exhibiting broad applicability and robustness against unlearning attacks.